OpenStudy (anonymous):

Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 − 2x − 7 10x2 − 7x − 3 8x2 − 11x + 3

OpenStudy (anonymous):

2x2+13x+15=(2x+3)(x+5) 2x2+27x+13=(2x+1)(x+13) 5x2-2x-7=(5x-7)(x+1) 10x2-7x-3=(10x+3)(x-1) 8x2-11x+3=(8x-3)(x-1)

OpenStudy (anonymous):

choose 2